OpenStudy (anonymous):

Hey guys, I have another question, this one involves the quadratic formula. Trying to find all solutions for x^2-5=squareroot of 3x using the quadratic formula

OpenStudy (anonymous):

\[x ^{2}-5=\sqrt{3x}\]

OpenStudy (anonymous):

Was this an assigned problem? The quadratic formula isn't really applicable to this case.

OpenStudy (anonymous):

Solve the equation 2−5= root 3x using the quadratic formula. Approximate solutions to three decimals places.

OpenStudy (anonymous):

I know I would put it into A+B+C form like \[x ^{2}-\sqrt{3x}-5\] But when I try to plus it into the equation it turns into a mess

OpenStudy (anonymous):

You sure it's not supposed to be 2x-5 = sqrt(3x)

OpenStudy (anonymous):

sqrt(3x) = sqrt(3)*sqrt(x)

OpenStudy (anonymous):

or maybe \[\sqrt{3}x\]?

OpenStudy (anonymous):

I don't see how that helps him, Estudiar.

OpenStudy (anonymous):

I think you're miscopying the problem somehow. Either the x isn't supposed to be under the radical or that first term isn't x^2 but just x.

OpenStudy (anonymous):

because if you square both sides you are going to get a 4th degree polynomial with a linear term (degree 1) and no quadratic formula will help you there

OpenStudy (anonymous):

Yes. You could use the quartic formula! Hurrah!

OpenStudy (anonymous):

Ohhh yes! YOur right Smooth, it is \[\sqrt{3}\]x

OpenStudy (anonymous):

Thanks :p I make silly mistakes, I should be able to work it out now.

OpenStudy (anonymous):

Ah, well then that is simple. It becomes x^2 - sqrt(3)*x -5 = 0.

OpenStudy (anonymous):

Thanks for the help. I will go work it out.

OpenStudy (anonymous):

whew. now use \[x^2-\sqrt{3}x-5=0\] \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=1,b=-\sqrt{3},c=-5\]